OpenStudy (anonymous):

Hi. Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (haseeb96):

use this formula y-y1 = m(x-x1) since , m = -6 (x1,y1) = (-2,-3) y -(-3) = -6( x - (-2) ) y +3 = -6 ( x + 2) y + 3 = -6x -12 6x + y + 15 =0 is your answer

OpenStudy (anonymous):

Thanks! @Haseeb96